OpenStudy (anonymous):

Three consecutive odd integers are such that the square of the second is 96 less than the square of the third. Find the three integers. This hurts me!

myininaya (myininaya):

assume x is odd what is the next odd number after x x+1 is even so x+2 is odd next odd number is x+4 so we have (x+2)^2=(x+4)^2-96

myininaya (myininaya):

(x+2)^2=(x+2)(x+2)=x^2+2x+2x+4=x^2+4x+4 (x+4)^2=(x+4)(x+4)=x^2+4x+4x+16=x^2+8x+16 --------- x^2+4x+4=x^2+8x+16-96 x^2+4x+4=x^2+8x-80 x^2-x^2+4x-8x+4+80=0 0-4x+80=0 -4x=-80 x=20 x+2=22 x+4=24

OpenStudy (zarkon):

21,23,25

OpenStudy (zarkon):

I would let x be the 2nd number then solve \[x^2+96=(x+2)^2\]

myininaya (myininaya):

oops x^2+4x+4=x^2+8x-80 x^2-x^2+4x-8x=-80-4 -4x=-84 x=21 so x+2=23 and x+4=25

myininaya (myininaya):

i forgot to add 4 and 80 :(

myininaya (myininaya):

i think im dyslexic

OpenStudy (zarkon):

not good

myininaya (myininaya):

i don't think it ever helped anyone to be dyslexic

OpenStudy (zarkon):

no

OpenStudy (anonymous):

Awesome!